OpenStudy (josee):

A rock is thrown into a still pond and causes a circular ripple. If the radius of the ripple is increasing at a rate of 4 ft/sec, at what speed is the area of the ripple increasing when its radius is 6 feet?

OpenStudy (anonymous):

oh my favorite ridiculous word problem! i guess we have to do it.

OpenStudy (anonymous):

you know that \[A=\pi r^2\] and you are looking for \[\frac{dA}{dt}\] right? the change is A with respect to time.

OpenStudy (anonymous):

taking the derivative of \[A=\pi r^2\] we get \[A'(t)=2\pi r r'(t)\] by the chain rule. you are told that \[f'(t)=4\] so this means \[A'(t)=8\pi r\]

OpenStudy (anonymous):

sorry that should read \[r'(t)=4\]

OpenStudy (anonymous):

now replace r by 6 and you get your answer

OpenStudy (josee):

OHHHHH! Gotchaaaaaa